# FreeA: Human-object Interaction Detection using Free Annotation Labels

### v1

@@ IV Experiments
Explanation-HOI [53] ECCV 2020 - -
MAX-HOI [24] WACV 2021 - -
PPR-FCN [25] ICCV 2017 - -
Align-Former [44] arXiv 2021 15.82 16.34
PGBL [26] ICLR 2023 42.97 48.06
Ours - 50.25 52.05
@@
SCG* [54] ICCV 2021 20.05 -
Weakly-HOI [23] CVPR 2023 29.59 -
Weakly-HOI† [23] CVPR 2023 17.71 -
@@
- 30.82 32.60
Evaluation Metric. The mean Average Precision (mAP) is used as the evaluation metric. The correct HOI triplet prediction should satisfy: 1) the predicted human and object bounding boxes have an IoU (Intersection over Union) with the ground truth greater than 0.5, and 2) the predicted category is correct. Evaluation on the HICO-Det dataset encompasses three modes: Full (600 HOI categories), Rare (actions with fewer than 10 training instances), and Non-Rare (the rest HOI categories). For the V-COCO dataset. We assess two scenarios: S1 with 29 action categories (including 4 body actions) and S2 with 25 action categories.
Implementation Details. The learning rate of HOI model is 1e-4. The optimizer is AdamW with a weight decay rate of 1e-4. The encoder has 6 layers, and the decoder has 3 layers. The total number of training epochs is 90, with a learning rate drop of ten times after the 60th epoch. The $\omega_{1}$ in Eq. 12 and $\omega_{2}$ in Eq. III-D are set to 0.23 and 0.1, respectively. All experimental trials are executed on 2 NVIDIA A800 (80G) GPUs, employing a batch size of 16. The computational environment runs Ubuntu 22.04, with Python version 3.8, PyTorch version 1.7.1, torchvision version 0.8.2, and CUDA version 11.0.
IV-A Effectiveness for Regular HOI Detection
As illustrated in Table I, we conduct comparisons under different supervised scenarios using the HICO-DET dataset. In contrast to the state-of-the-art (SOTA) weakly supervised Weakly-HOI method with image-level $\langle$“interaction”$\rangle$ labels for training, our method demonstrates a relative 8.58 mAP improvement in the absence of manual labels. Furthermore, we extend our method to weakly+ supervised field. Compared with the SOTA weakly+ supervised PGBL method, our method achieves better performance in both full and non-rare scenarios under default setting, with a performance improvement of 1.68 mAP and 2.48 mAP, respectively. It is noteworthy that our approach surpasses several fully supervised HOI models (e.g., InteractNet, iCAN, IDN). As shown in Figure 5, the proposed FreeA can almost generate comparable HOI labels with ground truth.
@@ IV-A Effectiveness for Regular HOI Detection
Figure 5: Comparison of HOI labels. GL represents labels generated by our method, and GT represents ground truth. In each sub image, the red and blue rectangles are bounding boxes for the human and object, and the green lines represent the connection between their centers. Green text indicates correct interactions.
IV-B Ablation Studies
IV-B1 HOI model
To evaluate the effectiveness of HOI model to the FreeA, four up-to-date fully supervised HOI approaches are tested, as shown in Table III. It is observed that a better HOI model can promote the overall effect of FreeA. For example, the QPIC model, with a 29.07 mAP in terms of full under fully supervised, achieves a 20.18 mAP in the weakly+ supervised. When we employ a superior HOI model, such as GEN-VLKT, it reaches 24.57 mAP in the weakly+ supervised. In this work, the GEN-VLKT model is applied as the HOI baseline model.
TABLE III: Ablation study using different HOI models in weakly+ supervised on HICO-Det datasets. $\tau$ represents performance under full supervision.
Method Source Defalut (mAPnormal-$\uparrow$)
@@ IV-B1 HOI model
[…]
IV-B2 HOI text templates
We conducted additional ablation studies to investigate various components of in the FreeA framework, and the results have been tabulated in Table IV. The experimental results show that both HOI text templates $T_{1}$ and $T_{2}$ play a vital role in HOI detection, at a 1.14 mAP increase as compared to FreeA with $T_{1}$ (Row 1 and Row 2 in Table IV). This is mainly because we observe that $T_{1}$ text template (“a photo of a person verb-ing an object) does not capture significant differences in similarity between different actions on the same object. Therefore, we introduce the $T_{2}$ text template, which emphasizes the action (“a photo of a person verb-ing”).
TABLE IV: Ablation study using different modules on HICO-Det datasets.
@@ IV-B2 HOI text templates
(delete background)
ICM
mAPnormal-$\uparrow$
1 $\checkmark$ $\checkmark$ $\checkmark$ 12.87
2 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 14.01
3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 14.87
4 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 15.83
5 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 16.14
6 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 16.96
Figure 6: HOI detection samples in HICO-Det. We retain only the human-object interaction pairs with the highest probability of interaction in the images. The first two rows showcase samples of HOI for different actions. The third row displays HOI samples for the same action with different objects. The fourth row illustrates HOI samples for different actions with the same object. The last row demonstrates that our method can also identify examples with no interactions.
Figure 7: HOI detection samples in V-COCO. The first two rows show HOI samples for different actions involving different objects. The third and fourth rows depict HOI samples for the same action with different objects and different actions with the same object, respectively.
IV-B3 Action selection approaches
Two action selection approaches, namely, “$top$-1” and “adaption”, are designed to determine which action should be selected when human-object interactions are not present in an image (Row 2 and Row 3 in Table IV). The “$top$-1” refers to selecting the most salient action, whereas “adaption” (Eq. III-D) retains actions within a specified threshold range. The results show that the “adaption” approach outperforms the “$top$-1” approach at +0.86 mAP. The $top$-1 only selects the action with the highest similarity, however, HOI typically involve interactions with multiple actions, and the “adaptation” method can satisfy this to provide multiple choices.
IV-B4 Dynamic threshold
We further test the effect of dynamic threshold $\theta$ (Eq. 12) to the FreeA (Row 3 and Row 4 in Table IV). It is observed that using dynamic threshold instead of fixed threshold results in a 0.96 mAP improvement. This indicates that the variability in the subtraction of the average similarity of all actions from the highest similarity action obtained through CLIP is significant, and fixed threshold cannot overcome this problem.
IV-B5 Background retention or deletion
The image background can be a double-edged sword. Sometimes it works in your favour when for simple image background, sometimes it works against you when one includes complex background details leading to different interferences. We conducted experiments to verify the effect of image background. The results indicate that retaining the background leads to performance decrease (Row 4 and Row 5 in Table IV).
IV-B6 The ratio of background
The effects of retention and deletion background information are presented in Table V. The ratio of 1.0:1.0 represents feeding two types of images separately into CLIP and then combine the final similarity results in a 1:1 ratio. Through experiments, it is found that as the proportion of deletion background image increased, the mAP results of FreeA also increased continuously. The last row in Table V indicates using only deletion background image, and it can be observed that in this case, FreeA achieves the the best performance.
TABLE V: Ablation studies using different segmentation images on HICO-Det datasets. The “()” represents the sum ratio of the results. For example, “1.0:1.0” means that two types of images are input into CLIP, and the results are combined in a 1:1 ratio.
Segmentation
(retrain
background)
Segmentation
(delete
background)
mAPnormal-$\uparrow$
$\checkmark$(1.0) $\checkmark$(1.0) 16.07
$\checkmark$(0.5) $\checkmark$(1.5) 16.37
$\checkmark$(0.25) $\checkmark$(1.75) 16.41
$\times$(0.0) $\checkmark$(2.0) 16.96
IV-B7 ICM
Regarding the proposed ICM component, the experimental results demonstrate an improvement of 0.82 mAP when ICM is utilized compared to when it is not (Row 5 and Row 6 in Table IV). The improvement is foreseeable because the ICM module emphasizes the correlation between actions.
TABLE VI: Ablation study using different modules on V-COCO datasets.
Row $T_{1}$ $T_{2}$ $top$-1 Adaption
Dynamic
threshold
Segmentation
(retain background)
Segmentation
(delete background)
ICM
mAPnormal-$\uparrow$
@@ IV-B7 ICM
1 $\checkmark$ $\checkmark$ $\checkmark$ 25.06 (23.77)
2 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 25.14 (24.67)
3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 27.01 (28.18)
4 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 27.30 (28.19)
5 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 30.11 (31.55)
6 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 30.82 (32.60)
IV-B8 The effectiveness for various componets of FreeA in V-COCO
We also conduct ablation experiments of each component of FreeA on the V-COCO, as shown in Table VI. Similar to the results on the HICO-Det, on the V-COCO dataset, using two text templates ($T_{1}$ and $T_{2}$) performs better than using only the $T_{1}$ text template (Row 1 and Row 2), and the “adaption” strategy is more effective than “$top$-1” (Row 2 and Row 3). The dynamic threshold method is also proven effective on the V-COCO (Row 3 and Row 4). Regarding the issue of whether to retain the background, the results on the V-COCO also indicate better performance when deleting background information (Row 4 and Row 5). Finally, by introducing the ICM module, the experimental results are further improved (Row 5 and Row 6).
V Visualization
@@
We visualize some results of generated labels compared with ground truth HOI labels, as shown in Figure 5. Most of the time the generated labels match the ground truth labels perfectly (row1 and row2).
We visualize the proposed FreeA method on two datasets, HICO-det and V-COCO, without using any manually annotated labels. For clarity, we retain only the human-object interaction pairs with the highest probability of interaction in the images. Figure 6 shows the visualizations of FreeA on HICO-Det, and Figure 7 presents the visualizations of FreeA on V-COCO. As shown in Figure 6, the first two rows display the detection results of FreeA on different actions. The third row lists cases where the same action is performed with different objects. The fourth row visualizes different actions involving the same object, and the last row illustrates the detection performance of FreeA in the absence of interacting human-object pairs. This demonstrates that FreeA can effectively detect various types of interactions.
For the V-COCO, the first two rows of Figure 7 visualize the detection results for different actions. The third and fourth rows show cases of the same action with different objects and the same object with different actions, respectively.
VI Conclusion
@@
We propose a novel weakly-- supervised HOI detection method, termed as FreeA. Weakly-- supervised FreeA means the training labels are not manually annotated from the raw datasets, but automatically generated from CLIP with the combination of candidate image construction, human-object potential interaction mining and human-object interaction inference modules. As compared to those weakly, weakly+, and fully supervised HOI methods, extensive experiments have demonstrated the effectiveness and advantages of the proposed FreeA. Our contributions to the field include presenting a new problem of weakly supervised HOI detection, and showing the utilization of CLIP model for generating HOI labels.

### v2

@@ 4 Experiments
Align-Former ResNet-101 15.82 16.34
Weakly HOI-CLIP ResNet-50 42.97 48.06
OpenCat ResNet-101 34.40 36.10
Ours ResNet-50 50.25 52.05
Ours Swin-L 57.66 59.78
@@
SCG* ResNet-50 20.05 -
VLHOI ResNet-50 29.59 -
VLHOI† ResNet-50 17.71 -
@@
ResNet-50 30.82 32.60
Ours (no labels)
Swin-L 35.01 37.30
Table 3: Performance comparisons on V-COCO dataset with different object detectors and backbones.
@@
weakly+ supervised
Weakly HOI-CLIP Faster R-CNN ResNet-50 42.97
OpenCat Faster R-CNN ResNet-101 34.40
Ours Faster R-CNN ResNet-50 46.32
Ours DETR ResNet-50 48.11
Ours Yolov8 ResNet-50 50.25
Ours Yolov8 Swin-L 57.66
weakly supervised
VLHOI† Faster R-CNN ResNet-50 17.71
Ours (no labels)
Faster R-CNN ResNet-50 28.12
Ours (no labels)
DETR ResNet-50 29.93
Ours (no labels)
Yolov8 ResNet-50 30.82
Ours (no labels)
Yolov8 Swin-L 35.01
4.1 Effectiveness for Regular HOI Detection
As shown in Table 1, we compare our method with various supervision levels on HICO-DET. Under weakly supervision using only “interaction” labels, our ResNet-50-based approach surpasses the SOTA VLHOI method by 7.47 mAP—without requiring manual annotations. With a Swin-L backbone and a YOLOv8 detector, our method achieves 21.67 mAP in the Full setting. Extending to the weakly+ setting, we exceed the latest Weakly HOI-CLIP method by 1.25 and 1.06 mAP in the Full and Non-Rare settings, respectively, using ResNet-50 and Faster R-CNN. However, performance on Rare categories is lower due to our pseudo-labeling strategy: when passed through CLIP, rare-class actions often yield low similarity scores and are thus filtered out, impacting recognition accuracy. Additionally, with ResNet-101 and Faster R-CNN, our method outperforms OpenCat across all metrics. Using a Swin-L backbone, we surpass OpenCat by 7.19, 7.78, and 7.08 mAP in the Full, Rare, and Non-Rare settings, respectively. Notably, our method also outperforms several fully supervised HOI models (e.g., InteractNet, iCAN, QPIC, RCL).
[…]
Table 5 presents the results of our proposed method under the zero-shot learning setting. RF-UC denotes the rare first setting, while NF-UC refers to the non-rare first unseen setting. UC and UV represent the unseen composition and unseen verb settings, respectively [22, 33]. In HOI, zero-shot learning involves training the model on a subset of seen interaction categories and testing it on previously unseen categories. Our method directly generates predictions for the unseen categories and then trains the HOI model, thereby enabling evaluation under zero-shot learning conditions. As shown in the table, our method achieves promising results across various zero-shot settings, demonstrating its effectiveness and generalization capability in handling previously unseen interaction combinations.
4.2 Ablation Studies
HOI model. Our proposed method is plug-and-play. To evaluate the effectiveness of HOI model to the FreeA, four up-to-date fully supervised HOI approaches are tested, as shown in Table 5. It is observed that a better HOI model can promote the overall effect of FreeA.
For example, the QPIC model, with a 29.07 mAP in terms of full under fully supervised, achieves a 20.18 mAP in the weakly+ supervised. When we employ a superior HOI model, such as GEN-VLKT, it reaches 24.57 mAP in the weakly+ supervised. After replacing it with the more advanced RLIPV2 model, the results were further improved, reaching 33.01 mAP, 32.13 mAP, and 33.27 mAP, respectively. To conduct ablation experiments efficiently, the upcoming tests will use GEN-VLKT as our HOI model.
Table 4: Performance comparison for zero-shot HOI detection on HICO-Det.
[…]
Full Unseen Seen
HOICLIP [22] UC 32.99 34.85 25.53
KI2HOI [33] UC 34.56 35.76 27.43
Ours UC 34.91 35.84 28.14
HOICLIP [22] UV 31.09 32.19 24.30
KI2HOI [33] UV 31.85 32.95 25.20
Ours UV 33.39 35.22 26.47
HOICLIP [22] NF-UC 27.75 28.10 26.39
KI2HOI [33] NF-UC 27.77 28.31 28.89
Ours NF-UC 29.15 29.93 29.34
HOICLIP [22] RF-UC 32.99 34.85 25.53
KI2HOI [33] RF-UC 34.10 35.79 26.33
Ours RF-UC 34.86 35.19 28.04
Table 5: Ablation study using different HOI models in weakly+ supervised on HICO-Det datasets. $\tau$ represents performance under full supervision.
[…]
@@ 4.2 Ablation Studies
35.10τ
RLIPv2
ICCV
33.01
43.23τ
32.13
45.64τ
33.27
45.09τ
HOI text templates. We conducted additional ablation studies to investigate various components of FreeA, and the results have been tabulated in Table 6. It shows that both HOI text templates $T_{1}$ and $T_{2}$ play a vital role in HOI detection, at a 0.93 mAP increase as compared to FreeA with $T_{1}$ (Rows 1 and 2). This is mainly because we observe that $T_{1}$ text template (“a photo of a person verb-ing an object) does not capture significant differences in similarity between different actions on the same object. Therefore, we introduce the $T_{2}$ text template, which emphasizes the action (“a photo of a person verb-ing").
Figure 4: Comparison of HOI labels. GL and GT represents generated labels and ground truth, respectively. The red and blue rectangles are bounding boxes for the human and object, and the green lines represent the connection between their centers. Green text indicates correct interactions.
Table 6: Ablation study using different modules on HICO-Det datasets using Swin-L backbone and Yolov8 detector under weakly+ settings. RB denotes retaining the background. DB indicates deleting the background.
@@
ICM PKM
@@
1 $\checkmark$ $\checkmark$ $\checkmark$ 25.83
2 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 26.76
3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 27.85
4 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 29.03
5 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 29.94
6 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 31.48
7 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 33.01
Action selection approaches. Two action selection approaches, namely, “$top$-1" and “adaption", are designed to determine which action should be selected when human-object interactions are not present in an image (Row 2 and Row 3 in Table 6). The “$top$-1" refers to selecting the most salient action, whereas “adaption" (Eq. 11) retains actions within a specified threshold range. The results show that the “adaption” approach outperforms the “$top$-1" approach at +1.09 mAP. The $top$-1 only selects the action with the highest similarity, however, HOI typically involves interactions with multiple actions, and the “adaptation" method can satisfy this to provide multiple choices.
Dynamic threshold. We further test the effect of dynamic threshold $\theta$ (Eq. 10) on the FreeA (Row 3 and Row 4 in Table 6). It is observed that using a dynamic threshold instead of a fixed threshold results in a 1.18 mAP improvement. This indicates that the variability in the subtraction of the average similarity of all actions from the highest similarity action obtained through text-image matching model is significant, and a fixed threshold cannot overcome this problem.
Background retention or deletion. The image background can be a double-edged sword. Sometimes it works in your favor for simple image background, sometimes it works against you when one includes complex background details leading to different interferences. We conducted experiments to verify the effect of image background. The results indicate that retaining the background leads to a performance decrease (Row 4 and Row 5 in Table 6).
ICM and FCM. Regarding the proposed ICM component, the experimental results demonstrate an improvement of 1.54 mAP when ICM is utilized compared to when it is not (Row 5 and Row 6 in Table 6). The improvement is foreseeable because the ICM module emphasizes the correlation between actions. Furthermore, comparing rows 6 and 7 in Table 6, we observe an additional gain of 1.53 mAP after incorporating the PKM module, demonstrating its effectiveness in enhancing performance by leveraging prior knowledge of object-action associations.
5 Visualization
@@
We visualize some results of generated labels compared with ground truth HOI labels, as shown in Figure 4. It’s clear that the generated labels have a high overlap with the ground truth labels.
6 Conclusion
@@
We propose a novel weakly-- supervised HOI detection method, termed FreeA. Weakly-- supervised FreeA means the training labels are not manually annotated from the raw datasets, but automatically generated from the text-image matching model with the combination of candidate image construction, human-object potential interaction mining, and human-object interaction inference modules. Compared with those weakly, weakly+, and fully supervised HOI methods, extensive experiments have demonstrated the effectiveness and advantages of the proposed FreeA. Our contributions to the field include presenting a new problem of weakly supervised HOI detection and showing the utilization of the text-image model for generating HOI labels.